### v1

# LeapFrog: The Rowhammer Instruction Skip Attack

Abstract
Since its inception, Rowhammer exploits have rapidly evolved into increasingly sophisticated threats not only compromising data integrity but also the control flow integrity of victim processes. Nevertheless, it remains a challenge for an attacker to identify vulnerable targets (i.e., Rowhammer gadgets), understand the outcome of the attempted fault, and formulate an attack that yields useful results.
In this paper, we present a new type of Rowhammer gadget, called a LeapFrog gadget, which, when present in the victim code, allows an adversary to subvert code execution to bypass a critical piece of code (e.g., authentication check logic, encryption rounds, padding in security protocols). The LeapFrog gadget manifests when the victim code stores the Program Counter (PC) value in the user or kernel stack (e.g., a return address during a function call) which, when tampered with, re-positions the return address to a location that bypasses a security-critical code pattern.
This research also presents a systematic process to identify LeapFrog gadgets. This methodology enables the automated detection of susceptible targets and the determination of optimal attack parameters. We first showcase this new attack vector through a practical demonstration on a TLS handshake client/server scenario, successfully inducing an instruction skip in a client application. We then demonstrate the attack on real-world code found in the wild, implementing an attack on OpenSSL. Our findings extend the impact of Rowhammer attacks on control flow and contribute to the development of more robust defenses against these increasingly sophisticated threats.
1 Introduction
The miniaturization of DRAM technology has inadvertently increased the susceptibility to bit flips and reliability issues. To mitigate data corruption, DRAM rows are refreshed at regular intervals, typically every 64 milliseconds. However, Kim et al. [25] discovered that rapid and repeated access to adjacent rows could accelerate charge leakage, leading to bit flips prior to the scheduled refresh, a phenomenon known as the Rowhammer effect [25]. Expanding on this, Seaborn et al. [43] demonstrated an even more efficient method known as double-sided Rowhammer, which exacerbates the issue.
Further developments in exploiting the Rowhammer vulnerability have been numerous. Gruss et al. [19] achieved root privileges by flipping opcodes in the sudo binary through a single-location hammering technique. Additionally, the feasibility of launching Rowhammer attacks remotely via JavaScript was shown by Gruss et al. [20] and Ridder et al. [17]. Tatar et al. [45] and Lip et al. [31] extended the reach of Rowhammer to network-based attacks. Its applicability in cloud environments [51, 13] and on FPGA-CPU hybrid platforms [50] has also been demonstrated. Importantly, Kwong et al. [29] revealed that Rowhammer poses not only an integrity threat but also compromises confidentiality.
Efforts to detect  [24, 12, 55, 23, 37, 21, 6, 16] and neutralize [20, 47, 8] Rowhammer attacks have been substantial. Yet, Gruss et al. [19] demonstrated the ineffectiveness of these countermeasures. Moreover, the security of ECC as a countermeasure was questioned by Cojocar et al. [14]. The Target Row Refresh (TRR) hardware countermeasure was recently circumvented as well, as shown by Frigo et al. [18] and further exploited by Ridder et al. [17] to target DDR4 chips with TRR. Notably, a recent study by Kogler et al. [27] highlighted the feasibility of hammering beyond adjacent locations to bypass TRR defenses.
Software exploits enabled by Rowhammer were studied further in recent studies. Tobah et al. [1] introduced the notion of Rowhammer gadgets and a specialized attack. Specifically, if a victim code is designed to return benign data to an unprivileged user, and uses nested pointer dereferences, Rowhammer can be used to flip these pointers thereby gaining arbitrary read access in the victim’s address space. Adiletta et al. [3] demonstrated that even CPU internal elements like register values, which are occasionally saved to the stack, can be vulnerable to Rowhammer when they are temporarily stored in the stack and flushed to memory. Upon reloading, these corrupted values are returned to the registers, potentially leading to execution of faulty stack variables and security breaches.
The threat of physical fault injection attacks has been acknowledged in the cryptographic community for some time [7]. For instance, OpenSSL incorporated error checks in CRT-based exponentiation early on to combat Bellcore attacks [7]. However, fault injection techniques have successfully compromised Elliptic Curve Parameters in the OpenSSL library [44]. Similarly, a Rowhammer-induced fault in WolfSSL, leading to ECDSA key exposure, was revealed in [35]. The vulnerability occurred during the TLS handshake process, involving the signing operation with private ECC keys. WolfSSL responded by introducing WOLF_SSL_CHECK_SIG_FAULTS, a series of checks during the signing stages to detect data tampering [36].
A recent work by Adiletta et al. [3] targeted sensitive stack variables via Rowhammer threatening data integrity. In this paper we instead target the control-flow-integrity (CFI) and subvert the execution flow for malicious ends, e.g. to bypass sensitive sections of code user authentication and data encryption. For this we introduce LeapFrog a new Rowhammer attack vector that targets the PC when stored in the stack during function calls and context switches. Not all PC manipulations will yield useful results, as some jumps within the code will result in errors, like segfaults, or simply will not bypass the intended code logic. To explore the massive attack surface, we introduce an automated tool that dynamically analyzes code to detect this particular type of Rowhammer gadget [1].
1.1 Our contributions
We introduce a novel approach for identifying LeapFrog Rowhammer gadgets capable of corrupting the PC, utilizing a combination of GDB, the Intel Pintool, and the Linux Process Interface.
Our contributions are fourfold:
  1. 1.
    We introduce the concept of LeapFrog gadgets, which allows an attacker to bypass security critical areas of code by faulting the PC value stored in stack.
  2. 2.
    We validate the feasibility of this attack in practical scenarios by successfully bypassing a TLS handshake in standard OpenSSL implementations.
  3. 3.
    We introduce the first simulation tool designed to identify LeapFrog gadgets. This tool represents an improvement over existing methodologies [53] by systematically analyzing binaries with our Intel Pin-based tool called MFS and incorporating time-domain analysis in simulations.
  4. 4.
    We propose and evaluate countermeasures against the LeapFrog attack, offering insights into enhancing the resilience of systems against such advanced Rowhammer based exploits.
2 Background
Rowhammer DRAM is stored in an array architecture of memory cells, where a capacitor and a transistor form each cell in the array. Each cell, capable of storing a value of either 1 or 0, is connected along word lines that extend across the row. Additionally, bit lines intersect the word lines perpendicularly, linking them to each cell. When the bit lines are brought into opposition (where one goes high and one goes low), positive feedback from a sense amplifier sets the state of the cell to be high or low. The sense amplifier consists of two cross connected inverters between the cells. For the cell to retain its state, the sense amplifier must be disconnected [2].
The sense amplifiers must be disconnected to read the cell, so the target word-line must be brought high. The charge in the capacitor will bring one of the bit lines high if the cell has a value of 1 during reading, then the sense amplifier will be reconnected, and the row will have the sense amplifier outputs latched [39].
The Rowhammer attack works by abusing the sense amplifiers ability to set the state of the cell. The capacitor will leak voltage, and must undergo a refresh of voltage every 64 ms (or less) according to standard JETEC convention. Reading the cell introduces noise into the system due to fluctuating voltages, and the noise can be amplified by the sense amp. As semiconductor technology improves, transistors shrink in size and the operating voltage also shrinks. The resulting cells are then able to operate at a higher speed, and the noise margins are reduced. Shrinking in size increases the signal coupling across traces and devices and magnifies crosstalk. When combined with the lower operating voltage and sharper edges due to higher speeds, the ratio of the crosstalk to the supply voltage increases significantly as technology improves making it easier for an adversary to exploit this type of attack. These can result in errors being written to the DRAM [2].
Instruction Skipping Instruction-skipping attacks are a type of fault attack that targets the normal execution flow of a program, particularly in embedded systems and secure circuits. Historically these techniques often needed physical access due to the timing and precision required to skip instructions. For example, laser fault injection has been demonstrated in inducing instruction skips in AES (Advanced Encryption Standard) that result in leaking private encryption keys [10]. In another example [41], researchers demonstrate how electromagnetic fault injection can effectively induce instruction skipping in the ARMv7-M architecture, specifically targeting AES.
In terms of countermeasures for instruction skips specifically, [34] addresses the vulnerability of embedded processors to instruction skip attacks. The paper acknowledges that while countermeasures based on temporal redundancy have been proposed, they are not entirely effective against double fault injections over extended time intervals.
Kernel Stack vs User Stack The user stack operates in user mode and is operated by user-level processes. Each user process has its own stack that stores local variables, function parameters, return addresses, and the control flow of the program. This stack is limited in size and is specific to the user space process, ensuring isolation and security from other processes.
On the other hand, the kernel stack operates in kernel mode, a privileged mode of operation for the system’s kernel. Each thread of a process has its own kernel stack. This stack is used when the process executes system calls or when it is interrupted and the kernel needs to perform operations on behalf of the process. The kernel stack handles lower-level, system-critical operations like interrupt handling, system call implementation, and managing hardware interactions. It is kept separate from the user stack for security and stability, ensuring that user processes cannot directly access or interfere with kernel-level operations.
Data stored in the kernel stack includes CPU context for system calls, interrupt state information, and other kernel-specific data, while the user stack holds user-level process data like function calls and local variables. This separation reinforces the security and stability of the operating system by isolating user applications from the core kernel functions.
Program Counters Program Counters (PCs), also known as Instruction Pointers, hold the memory address of the next instruction to be executed by the CPU. This mechanism ensures that instructions are executed in the correct sequence. The value of the Program Counter is typically not stored in the stack; rather, it’s stored in a dedicated register within the CPU. During the execution of a program, the PC is automatically incremented after each instruction is fetched, pointing to the subsequent instruction. However, during certain operations like function calls and interrupts, the PC value may be changed abruptly to a new address. In such cases, the return address (the original PC value) is often stored in the stack to enable the program to return to the correct point in the program after the operation is complete. For user function calls, the PC value gets pushed to the user stack, but if there is an exception, signal handler, or system call, the PC gets pushed to the kernel stack. This mechanism facilitates the smooth flow of program execution.
Process Degradation Process degradation in computing refers to the intentional slowing down of a processor to create favorable conditions for certain types of attacks. A notable contribution in this field, HyperDegrade [4], combines previous approaches [5] with the use of simultaneous multithreading (SMT) architectures to significantly slow down processor performance, achieving a slowdown that is orders of magnitude greater than previous methods. It utilizes collateral Self Modifying Code (SMC) events to induce "machine clears", where the entire CPU pipeline is flushed, resulting in severe performance penalties. This process is triggered by cache line eviction, causing the invalidation of instructions in the victim’s L1 instruction cache, which the CPU may interpret as an SMC event. This mechanism amplifies the degradation effect, as instructions are sometimes fetched multiple times, leading to substantial slowdowns in CPU performance. This slowdown enhances the time granularity for FLUSH+RELOAD [22] attacks, enabling more effective exploitation of side-channel vulnerabilities in systems. The attack not only explores the implementation of this technique but also investigates the root causes of performance degradation, particularly focusing on cache eviction. Their findings have substantial implications in the realm of cryptography, as evidenced by the amplification of the Raccoon attack [33] on TLS-DH key exchanges and other protocols.
3 Related Work
Similar work [19] achieved privilege escalation through opcode flipping. Researchers loaded the sudo binary into memory from user space and flipped a bit in the opcode of the binary such that an incorrect password would cause authentication and the correct password would cause mis-authentication. They mention flipping conditional jumps that change program execution flow. While our work also can result in privilege escalation, we do so by attacking the process during run time and forcing the process to jump to an unexpected line of code by flipping the PC value.
To attack the binary during runtime, we had to overcome timing challenges as well as different detection problems to find vulnerable areas in the code. In [19] researchers used mmap to map the target binary into a vulnerable page in memory, demonstrating how memory waylaying and memory chasing techniques can force the mapped binary into the target page. This attack can potentially be mitigated by making the process execute only, and thus cannot be mapped with the mmap command. In contrast, our work can attack binaries that are unreadable from userland and are execute only. Additionally, our attack works on fundamentally different mechanics, so targets not susceptible to [19] may be susceptible to ours.
Another related work [1] demonstrates that code using nested pointer dereferences can corrupt bits in these pointers to reveal data to an unprivileged user. They demonstrate this vulnerability on ioctl given they can flood the kernel heap with data by spawning processes (a method they call "spraying"), increasing the probability a single bit-flip will point to malicious data in the heap that points to the location of secret data. Our work compliments and improves upon this prior work by increasing the number of vulnerable code patterns, since their work relies on the presence of specific code patterns that may not be present in the victim code.
Lastly, [54] demonstrates a Rowhammer attack methodology where researchers emulated Rowhammer bitflips on targets. They introduced the idea of simulating a flip in the EIP register value in the stack, which can force the execution to jump from kernel code to user code, like the ret2usr attack [28]. However, attacks that cause privilege escalation by jumping from kernel code to user code are mitigated by SMAP [15], which prevents the kernel from executing userland instructions. Our attack forces a process to jump within its own code space and privilege space, and thus is not affected by SMAP and introduces attack surfaces on new code patterns.
4 Threat Model
Similar to other Rowhammer attacks we assume the attacker is co-located on the same system as the victim  [25, 19, 51, 13, 20]. Co-location is a common threat model for many micro-architectural side-channel attacks and fault attacks  [30, 26, 11, 46, 48]. We do not assume root privilege or physical access to the machine. For some attacks, we assume that the attacker can send signals to pause it during the attack, such as sudo. However, we do not need to be root to send signals. Additionally, we assume that the system has TRR enabled, and bypass TRR with a multi-sided attack [18].
5 LeapFrog Attack
Figure 1: LeapFrog gadget in TLS handshake $addr_{src}$, the PC value that fault is injected into, is highlighted in blue. The new value after the fault injected, is highlighted in red. The fault is injected during the execution of the function call highlighted in green.
Central to our investigation is the concept of LeapFrog gadgets, an important element in the orchestration of Rowhammer-based attacks targeting the Program Counter (PC). These gadgets are exploitable in scenarios where a process undergoes a context switch or executes a function call, leading to the storage of the PC value in either the kernel or user stack. The ingenuity of LeapFrog gadgets lies in their susceptibility to Rowhammer-induced bit flips due to them being stored in DRAM, enabling an attacker to subtly alter the PC value. This manipulation is designed to redirect the execution flow to a different code segment, ideally with minimal bit changes due to the blunt nature of Rowhammer and the higher probability of finding a faulty memory location with few or one faulty bits in the right location. In this paper, we assume that we can successfully find 1 bit flip within a page that is in the right location to fault the PC value to force the intended instruction skip.
In Figure 1 the storage of the PC value occurs in the kernel stack during the execution of wait_receive. In this scenario, a malicious server can hold the client process at the wait_receive function while hammering the PC value to force the process to jump to a new location upon returning from the function. In our assembly code analysis in Figure 1, we observe the original PC value is an address 0x555555555478. Through strategic bit flips, this value can be altered to 0x555555555578, effectively enabling an instruction skip (skipping one or more instructions) and jumping from the function call in wait_receive directly to a later point in the execution, bypassing the critical server authentication check. The practicability of such attacks, however, hinges on the feasibility of achieving the desired bit flips, a central challenge to the effectiveness of LeapFrog gadgets in real-world scenarios. In this scenario, flipping the PC from 0x555555555478 to 0x555555555578 only require 1 bit flip, which is a reasonble assumption for Rowhammer.
However, tiny variations in the C code can change the resulting assembly code significantly. For example, consider the first approach for a TLS handshake, where the process allocates memory for a message to be signed. The C code and its corresponding assembly code are shown in Listing 1.
Listing 1: Combined C and Assembly code for original memory allocation
1// C Code
2unsigned char message[32] = "This is a message to be signed";
3int ret = send(client_fd, message, sizeof(message),0);
4
5// Assembly Code
60x0000555555555413: movdqa 0xce5(%rip),%xmm0
70x000055555555541b: mov \$0x20,%edx
Listing 2: Combined C and Assembly code for alternative memory allocation
1// C Code
2unsigned char *message;
3message = "This is a message to be signed";
4int ret = send(client_fd, message, sizeof(message),0);
5
6// Assembly Code
70x555555555413: lea 0xc6e(%rip),%r14
80x55555555541a: lea 0x60(%rsp),%r13
90x55555555541f: mov %r14,%rsi
Alternatively, using a different method to allocate memory for the message results in a variation in the assembly code. This alternative approach and its corresponding assembly code are presented in Listing 2.
In the code space, the alternative approach (Listing 2) takes 0x55555555541f - 0x555555555413 or 12 bytes of instructions, while the original approach (Listing 1) occupies 8 bytes of assembly instruction. Given the assumption that only one bit per page can be reliably flipped, identifying useful instruction skips that require a single bit change, as illustrated in Figure 2, is crucial. This example illustrates the challenge in manually inspecting source code to determine the impact of tiny variations on assembly instruction distances. Hence, profiling binaries becomes an important tool in this context.
Figure 2: The best LeapFrog gadgets require a single bit flip, where the distance between the two lines of code is a power of 2
5.1 Offline Memory Profiling
Finding Contiguous Memory
In this work, we employed DRAMA [38] to ascertain contiguous memory regions. DRAMA provides a direct and efficient method for reverse engineering the DRAM layout, facilitating the identification of contiguous memory. DRAMA exploits the physical structure of DRAM through a timing side channel to reveal addressing information, enabling us to determine contiguous memory regions and bank affiliations. This is crucial for Rowhammer, as the attack requires the targeted rows to be in adjacent physical locations within the same DRAM bank.
To optimize memory allocation, we define a structure that maps the DRAM and virtual addresses of each page, then allocate a large buffer to populate this structure with the corresponding addresses. By implementing a two-phase sorting process—initially sorting by DRAM bank, followed by sorting rows within each bank—we efficiently identify sequences of physically contiguous rows within the same bank. This streamlined approach eliminates redundancy and precisely targets contiguous memory regions, enhancing the overall memory management strategy.
After sorting, we traverse the array to find the longest sequence of contiguous rows within the same bank. This is achieved by comparing the DRAM row values, looking for sequences that are the same or increment by one. The result is a segment of memory, identified as a continuous chunk, which is critical for mounting a successful Rowhammer attack.
Executing the Rowhammer Bit Flip in a Multi-Sided Context Despite modern mitigation techniques against Rowhammer like Target Row Refresh (TRR), we are still able to induce flips in DDR4 memory by using a multi-sided [18] approach.
In the final phase of our attack, the task is to induce bit flips in the target memory location. This step marks the culmination of the profiling and memory manipulation processes. The challenge lies in the fact that while we can ascertain the occurrence of bit flips in a given row (a row that we deem "flippy"), pinpointing the exact memory bits affected after the attack is not straightforward. This is due to the inherent nature of Rowhammer, where the attacker does not possess direct control over the specific memory areas being altered.
However, the success of the attack is often evident through observable changes in the process’ state. For instance, a successful execution might manifest as an unauthorized bypass of security measures, or broken encryption output. This indirect outcome serves as a confirmation of the attack’s effectiveness. We further expand on this in §7.
6 Locating the PC in the Stack
To flip bits in the PC value with Address Space Layout Randomization (ASLR) enabled, the page that contains the PC value needs to be placed into the page with the bit that will flip during the Rowhammer attack. To do this, we use a method similar to that proposed in [40] where we deallocate a series of pages from the attacker process, launch the victim process, and experimentally determine some probability that the target data (in this case the PC) lands in the target location (the row with the flippy bits). We term the deallocation of pages "baiting" in this paper.
The profiling to determine the proper number of bait pages starts by allocating pages within the attacker’s process space, designated to be released as bait. The procedure involves releasing a substantial number of bait pages, recording their physical addresses, and then correlating these with the physical address of the target variable in the victim process. The number of pages consumed by the victim process before allocating the target variable was determined through this correlation.
In a recent work [3], the victim’s source code was altered to assign a unique value to the target register or stack variable, thereby making it identifiable in the memory during the profiling stage. This method is not possible with PC values as they are dependent on the compiler, so we introduce a new method to determine the number of bait pages required for the PC value.
The dynamic nature of the PC under ASLR implemented in the Linux kernel, necessitates a novel approach that involves identifying invariant values within the stack that serve as reliable fingerprints. These fingerprints are used to determine the PC’s offset relative to these constants, thereby facilitating the estimation of the required number of bait pages for effective targeting.
Figure 3: Finding constant values in the stack to create a fingerprint
6.1 Fingerprinting the Stack
As the PC’s address and value fluctuate with each process execution due to ASLR, our strategy leverages the relative stability of certain stack values and correlating an offset from those values. We first profile with ASLR disabled, knowing the target PC value in the stack from an assembly dump with GDB. We then determine an offset from the fingerprint as seen in Figure 4. Then with ASLR enabled, even with the PC value changing, the fingerprint remains identifiable and the offset from the fingerprint remains constant. The outcome is a refined understanding of the number of bait pages required to strategically position the PC, thus enhancing the precision of our Rowhammer attack in an ASLR-enabled environment. Fingerprinting only needs to be done once and is machine-independent.
Figure 4: Once the fingerprint is located, there is a constant offset from the fingerprint regardless of ASLR, and this can be used for bait page profiling for the eventual attack
The process begins by capturing snapshots of the stack at different instances and identifying unique values that persist across these snapshots. We implemented a Python script to automate this analysis. The script compares consecutive stack states, isolating values that remain unchanged— these become features of our fingerprints as seen in Figure 3. By calculating the address differences between these consistent values and tracking their occurrence across multiple iterations, we build a comprehensive profile of the stack’s layout. This profile is instrumental in pinpointing the location of the PC relative to the identified fingerprints and is versatile enough to be used on virtually any binary.
7 Automatic Detection of LeapFrog Gadgets with MFS
Figure 5: LeapFrog gadget detection using MFS framework
Based on how the LeapFrog gadgets occur in the binary described in §5, we developed a custom tool we call MFS (Multidimensional Fault Simulator) that relies on dynamic binary instrumentation and analysis. Since the attack happens on program counters and registers, which are invisible to high-level code, such as C/C++, it is not possible to do a static analysis of the source code. We put together a set of rules that enables us to collect, filter, and pinpoint the potential LeapFrog gadgets.
1
First, MFS collects the instruction traces, specifically, the address of instructions executed, for different inputs. For the purpose of detecting the gadgets that cause security exploits, MFS chooses critical input pairs that cause differences in the program’s control flow. Such inputs can be correct/incorrect private key pairs or passphrases for authentication programs. Together with the instruction addresses, we collect the execution time of each function executed. Since the return addresses of the functions with larger execution times will stay in the memory for a longer duration, they are potentially more viable targets.
2
MFS then computes the difference between two instruction traces to find the instruction addresses that are executed with correct input(s) but not executed with incorrect input(s). Note that this is an optional step to reduce the complexity of the following steps, and it comes with a cost of false negatives. Moreover, depending on the program and type of exploit, it may not always be possible to get multiple different traces; see §8.1. Alternatively, the whole instruction trace can be considered instead of only the difference.
3
MFS then looks for address pairs that hold the following conditions:
$d_{H}(\textrm{addr}_{\text{exec}}^{i},\textsf{addr}_{\text{return}}^{j})=1$ (1)
where $addr_{\text{exec}}^{i}$ is the address of the $i^{th}$ instruction that is executed, $addr_{\text{return}}^{j}$ is the return addresses of the $j^{th}$ call instruction, and $d_{H}$ is the Hamming distance between two addresses. $i$ and $j$ are bounded by the number of all instructions executed ($n$) and the number of call instructions executed ($m$), respectively. Although this operation has $O(m^{n})$ complexity, it can be implemented with bitwise xor and can be parallelized using multiple processor cores. The condition given in Equation (1) is determined by the Rowhammer fault model. Since multiple-bit flips on a precise target are much rarer and harder to control, MFS assumes we can only flip a single bit. Yet, the method is generic enough to cover other potential fault models, such as optical fault injection, where multiple-bit flips are more likely [9]. This step generates a list of pairs of addresses in the following format: $\{<addr_{src}^{k},addr_{dest}^{k}>\}$ where $addr_{src}^{k}$ is the $k^{th}$ instruction address that MFS targets in the binary’s execution with the input that we want to affect the control flow of, such as an incorrect private key, and $addr_{dest}^{k}$ is the corrupted instruction address after fault injection.
4
For each address pair we get from the list generated in the previous step, MFS starts a simulation session. MFS executes the binary again with the incorrect input and simulate a bit flip on the instruction address $addr_{src}$ to make it $addr_{dest}$. It is possible that certain instructions are executed multiple times in a single execution. To correctly cover that case in our fault model, we keep a counter variable for a specific instruction that increments every time the binary executes the same instruction. In a single execution of the original binary, if an instruction is executed NN times, we attempt the fault simulation $N+1$ times, until we no longer see the same instruction in the trace.
5
After the bit flip simulation, MFS continues the execution of the binary without further faults and observe the new behavior. The analysis of the new behavior is not a trivial task. There are several options where we can observe changes compared to the original execution. For instance, we can observe changes in the total number of executed instructions, the number of instructions that match with the correct input execution trace, the return code of the program, outputs to standard streams, ports that are accessed, functions calls, authentication result, etc. The choice of observable depends on the program under test. In this work, MFS uses the return codes, standard outputs/errors, and authentications on different case studies.
6
Once MFS has a list of PC values that potentially result in authentication or bypass with a single-bit flip, it then evaluates if they are practical to attack from a timing perspective. In some cases, a single bit flip will result in the desired behavior in a process but the attack window of time is too short to effectively attack the target. Additionally, the attack window needs to be long enough to allow for noise in the system - as processes will often take a variable amount of time to execute and get to the vulnerable area in the code where the PC value is shelved in the stack. MFS uses process degradation to increase the viability of LeapFrog gadgets, as slowing down a process artificially increases the attack window time. Note that this step is system-specific and it can be affected by the current processor/memory load. Although it is necessary to find viable targets in the list for an end-to-end attack, it does not guarantee that the other targets are not viable in different system configurations, or in different systems.
MFS starts the victim process and then immediately stops it with a SIGSTOP signal and it checks if the PC value is currently in the stack of the process. If not, the process is killed and restarted, and stopped after a slightly longer period of time, in a process we call time sweeping. The challenge is sending a SIGSTOP with the highest timing precision possible. Different implementations of signals will yield different timing resolutions. For example, Python has a signal library that can be used to generate signals similar to a bash script, but there is considerable delay and imprecision in the time it takes to send a signal.
7.1 Tool Implementation
We used Intel’s dynamic binary instrumentation framework, Pin, which allow for process analysis without altering its core behavior  [32] to implement 1 and 4 of MFS. Using Pin also makes it possible to find LeapFrog gadgets in binaries that do not have a source code since it does not require recompiling. In the context of MFS, Pin’s capabilities are harnessed to monitor the execution trace of a binary. This integration allows for a thorough analysis of potential LeapFrog gadgets by observing how changes in PC values influence program behavior. For each executed instruction, our tool outputs the virtual address of the instruction and disassembly of the machine code. If the instruction is a call instruction it also outputs the return address of the call, which is usually the PC value that is pushed onto the stack before executing the called routine. For every write to STDOUT and STDERR, the tool forwards a copy of the buffer to a text file for further analysis. To avoid the effect of overhead caused by instruction-based instrumentation, function timings are collected in a separate session on every function entry and exit.
2
is a simple comparison operation on the correct and incorrect execution traces implemented with diff command line tool in Linux.
3
is implemented in Python. MFS parses the instruction traces and computes the Hamming distance between the return addresses and instruction addresses of all executed instructions in the correct trace or the list of addresses we get from 2. The Hamming distances are calculated using the native bit_count function in Python followed by bitwise_xor in numpy library. The operation is parallelized on multiple cores to speed up the analysis.
The bit flip simulation part of MFS ( 4) is done using Pin which takes the address pairs and simulates every fault independently. The faults on PC values are implemented as direct jumps to the corrupted addresses by adding jmp $addr_{dest}$ after function returns. Since we add a direct jump to the target address by injecting a line of assembly with the the Pin tool, it is functionally equivalent to corrupting the PC value in memory.
5
filters the simulation results depending on the program and targeted exploit type. For different types of exploits, we filter by return code (§8.3), value in STDOUT (§8.1), and authentication result (§8.2).
6
The last stage of MFS takes the list of PC values generated from the previous steps and determines which are practical from a timing perspective. It does this by sweeping the process in the time domain determining when it needs to stop the process to find particular PC values in the stack.
We begin by defining when we want to start our sweep, and what interval we want to sweep at. For OpenSSL as an example, we started our sweep at 0ns and had an interval of 100ns. Generally, the higher the resolution of the sweep, the longer the simulation takes. However, a smaller interval increases the likelihood that we will successfully send a SIGSTOP at a time when the target PC value is in the stack.
In order to determine if a PC value is in the stack, we start the victim process as a non-root user on a sibling core to a core that we are attacking with SMC to degrade the performance. For example, in our tests we isolated cores 6 and 14 and triggered SMC events on core 14 while running the victim process on core 6. Once we have the process identification number (PID) of the process and send a SIGSTOP, we use the Linux process interface to check the stack for the PC value. We do this by looking at the /proc/[pid]/maps file to determine which offsets in the victim process’s address space contain the stack, and then we read from /proc/[pid]/mem at the offsets determined by /proc/[pid]/maps to find the PC values. The tool will generate a dictionary of stack addresses/values for the victim process that we can search through.
If during a sweep the tool finds the PC value in the stack, it will simulate a flip by overwriting that value with the new PC value determined by the previous steps to verify that the gadget does result in the intended behavior (privilege escalation, data leak, mis-authentication, etc…).
Generally, if MFS can successfully pass all stages of filtering with a particular LeapFrog gadget, we believe that it can be attacked and flipped with Rowhammer to cause the desired behavior.
8 Experiments
Experiment Setup The experiments are conducted on a system with Ubuntu 22.04.2 LTS with 6.2.0-37-generic Linux kernel installed. The system uses an Intel Core i9-9900K CPU with a Coffee Lake microarchitecture. We used a dynamic clock frequency rather than a static clock frequency to improve the practicality of the attack. End-to-end attack experiments are done on a single DIMM Corsair DDR4 DRAM chip with part number CMU64GX4M4C3200C16 and 16GB capacity. DRAM row refresh period is kept at 64ms, which is the default value in most systems. In all the experiments, we used 100s simulation timeout, since the fault simulations rarely cause infinite loops. We empirically observe that using the Python signals library, the target process could complete 34M cycles before the attacker can stop it, with a standard deviation of 2.7M cycles. Alternatively using a bash script, the victim process can only complete 18M cycles before it is stopped, with a standard deviation of 0.3M cycles. There is an order of magnitude difference in precision stopping a process with bash vs with Python.
Target Size #Inst.exec $d_{HD}$ # Candidates
2 on 2 off
TLS 29KB 5328007 1 315 2493
2 2240 14413
3 21841 67421
OpenSSL 818KB 49431 1 N/A 2700
2 N/A 20208
3 N/A 70475
sudo 227KB 148177 1 1100 8655
2 6910 54203
3 30908 221181
Table 1: Number of gadget candidates found by MFS in for fault models with different Hamming distances. We ran OpenSSL with aria-128-cbc cipher.
8.1 OpenSSL Encryption Bypass
We analyze openssl command line tool that uses OpenSSL v1.1.1w for block cipher and stream cipher implementations. For each cipher, we give a simple plaintext that contains the helloworld string and run encryption without salt with a simple passphrase. Our aim is to find LeapFrog gadgets in the binary that can be exploited for bypassing encryption steps in the ciphers, revealing the plaintext.
First, we scan the binary using MFS as described in §7. Since we do not aim for any authentication bypass in this scenario, and the execution traces are deterministic for fixed inputs, step 2 is not applicable. Instead, in step 3, we compare the return addresses in a single trace against all the instruction addresses in the same trace to look for targets with $d_{HD}=1$.
We scanned the binary with 135 different ciphers available in OpenSSL. Most of the time the binary was not affected by the simulated bit flip and correctly produced the ciphertext.
Fig. 8 illustrates one of the LeapFrog gadgets found in the openssl command line tool. When we corrupt a single bit in 0x55555559c4c5, the return address of opt_cipher function, to make it 0x55555559c0d5, the function returns to the corrupted return address, skipping three instructions in between. Similarly, another single-bit corruption to (0x55555559c0c5) causes the function to return to an earlier point in the program. We verified that both of these bit flips cause the binary to skip the whole encryption and instead output the plaintext. Similarly, MFS detected LeapFrog gadgets that are used in 36 ciphers including block ciphers and stream ciphers. The ciphers with LeapFrog gadgets that revealed full or partial plaintext are listed in Table 2. Fig. 6 and  7 summarize the simulation results for aes-256-ctr and aria-256-ctr respectively.
Even with ASLR enabled, these gadgets are reproducible because ASLR does not randomize the last 12 bits of the code space (the page offset). We only simulated faults in the last 12 bits (which should be the same across all x86 machines the process is compiled for), thus, the LeapFrog gadgets should work across machines without the need for rescanning.
Figure 6: aes-256-ctr simulation results
Figure 7: aria-256-ctr simulation results. Plaintext helloworld is revealed three times.
Recovered Cipher
helloworld aria-128-cbc, aria-128-cfb,aria-128-cfb1
aria-128-cfb8, aria-128-ctr, aria-128-ofb
aria192, aria-192-cbc, aria-192-cfb
aria-192-cfb1, aria-192-cfb8, aria-192-ctr
aria-192-ofb, aria256, aria-256-cbc
aria-256-cfb, aria-256-cfb1, aria-256-cfb8
aria-256-ctr, aria-256-ofb, bf-ofb
rc2-ofb, rc4, rc4-40
hellowor bf-cfb, rc2-cfb
hdlmowor idea-cfb, idea-ofb
oworhell bf, bf-cbc, bf-ecb, blowfish
?rl#a?gy?... chacha20, des-ede3-ofb, des-ede-ofb, des-ofb
Table 2: 36 ciphers implemented in OpenSSL that are vulnerable to LeapFrog attack. Each given cipher reveals the plaintext fully or partially in the ciphertext due to skipped encryption steps.
[…]
8.2 Sudo Privilege Escalation
We analyze the sudo binary installed on the system for potential LeapFrog gadgets. Our objective is to bypass native password authentication implementation in sudo v1.9.9 resulting in privilege escalation.
First, we create a dummy C program that tries to bind to port 80 in the system, which is a privileged operation. As an unprivileged user, when we try to execute the program it outputs Bind failed: Permission denied error and exits. When we run the same program with root privileges, it outputs Successfully bound to port 80 and exits. The purpose of creating such a dummy program is to have a probe on the simulation that detects the attack’s success.
Then we run the program with sudo twice, first with the correct password and second with an incorrect password. This results in two different traces. We first run our detection tool with step 2 on, where we restrict ourselves to only a subset of instructions. With this mode, our tool we found 1100 unique gadget candidates. However, after running the rest of the filtering steps, none of them were found to be an actual working gadget.
Next, we turned off step 2 to increase the coverage. Our detection tool found 8655 unique candidates with $d_{HD}=1$ within the sudo binary and other dynamically linked libraries combined. After simulating these candidates, we found that 10 unique $addr_{src},addr_{dest}>$ pairs result in privilege escalation. We illustrate three of these address pairs in Fig. 9. Interestingly, $addr_{src}$ in those three gadgets are the same PC value (0x555555579970) which is the return address of the <sudo_strtoid_v2@plt> function. This shows that sudo will return to a point that allows unprivileged users to run as root with an incorrect password in case the malicious user flips one of the three identified bits during the execution of the said function. The remaining 7 gadgets we identified appear during the call instructions to <strncmp@plt> (5 times), <sudo_parse_gids_v1@plt> (1 time, also visible in Fig. 9), and <sudo_debug_enter_v1@plt> (1 time) functions. Note that stncmp function is implemented in libc library. This is a real-world example of a LeapFrog gadget caused by a combination of the instruction order in the main program binary and the PC storage in the stack during a third-party function execution. Therefore, LeapFrog mitigations should consider the role of library dependencies as well. Since a potential mitigation should be generic and independent of the program logic, the analysis of how these specific corruptions cause privilege escalation is not relevant and we leave it outside of the scope of this work.
Figure 9: LeapFrog gadgets detected in sudo binary. The PC value that fault is injected into, $addr_{src}$, is highlighted in blue. The new value after the fault injected, $addr_{dest}$, is highlighted in red. The fault is injected during the execution of the function call highlighted in green.
8.3 TLS Handshake
Figure 10: Probability distribution of bait page numbers.
In a full end-to-end attack example, we illustrate the potency of the attack by applying it within a client/server authentication framework, specifically using OpenSSL for signature verification. Here, we consider a scenario where the attacker shares a physical compute space with the client. The goal of the attacker is to manipulate the client’s signature verification mechanism, causing it to erroneously validate a corrupted signature as genuine. This manipulation forms part of a broader man-in-the-middle strategy, aimed at deceiving the client into believing they are securely connected to the intended server.
In the standard communication flow, the client initiates contact with the server by dispatching a ClientHello message. The server replies with a ServerHello message, which carries its public key and a digital signature of the handshake process. The client’s role is then to authenticate this signature using the server’s public key. Under normal circumstances, a verified signature would indicate a secure channel, prompting the client to transmit sensitive data to the server. However, in our attack scenario, the attacker strategically alters the signature verification process at the client’s end. By inducing a single-bit error during this process, the client is misled into accepting a fraudulent signature as valid. As a result, the client erroneously trusts the communication channel and proceeds to send sensitive information to the attacker.
Figure 11: TLS Handshake: The client attempts to authenticate the server, and a colocated rowhammer attacker flips the PC value causing an instruction skip resulting in a misauthentication - this is an end-to-end attack
Figure 11 illustrates a standard interaction where the client establishes a connection with the server, sends a request, and then receives a server-signed message, enabling server authentication. A critical aspect to note is the client’s susceptibility to a Rowhammer attack while it awaits the server’s response. This waiting period, which can last several milliseconds, is primarily dictated by the server’s response time. During this interval, an attacker has the opportunity to exploit the Rowhammer vulnerability by targeting the client’s memory.
We scan the client binary using MFS while the server is using the correct and incorrect private key. With step 2 on, we found 315 unique gadget candidates with $d_{H}=1$. When the server uses the correct key, the client binary terminates with return code 0, and when the server uses an incorrect key, the client returns with code 1. In step 5, we look for PC corruptions that cause the client to return with value 0, meaning it incorrectly authenticates the server. After simulation, we found that one of the candidates was a LeapFrog gadget that caused false authentication of the malicious server with an incorrect key.
Then, we scan the client with step 2 off. With this mode, MFS detected 2493 gadget candidates with $d_{H}=1$. After the simulation steps, we verified that 21 of those candidates were LeapFrog gadgets that cause the client to return with 0, including the one found earlier. The number of candidates for different Hamming distance values are given in Table 1.
The total time for the end-to-end attack to induce a successful misauthentication of the TLS handshake was 12 hours and 25 minutes, as seen in Table  3. This time included profiling the system for the proper flippy pages with the correct offset, meaning the actual online time was around 2 hours. The experiment found a total of 1647 unique flippy pages, and over the course of the 2 hours of online attacking, we saw 2206 attacks where the program counter was baited into the correct page we were attacking before it flipped.
Category Result
Total Time 12 hrs 25 mins
Online Time 1 hr 54 mins
Total Flippy Pages 1647
Total Attacks w/ Correct # of Bait pages 2206
Table 3: Results from the end-to-end attack on code using OpenSSL client/server signature verification with LeapFrog gadget
9 Countermeasures
9.1 Rowhammer Resistant Hardware
Increasing the DRAM refresh rate is a commonly cited countermeasure to prevent Rowhammer attacks. Standard DRAM refresh is 64ms, meaning that a Rowhammer attack has 64ms to flip a bit before the row refreshes. Thus a faster refresh rate will result in a shorter time window for the Rowhammer attack to be performed and should result in fewer flips. This is not an ideal solution however, because a faster refresh rate will lead to worse power usage and performance overall. Alternative methods such as probabilistic row refresh [49] and parallel row refresh [52] are not available in consumer systems.
A novel countermeasure against Rowhammer attacks is the Randomized Row-Swap (RRS) method [42]. This approach fundamentally disrupts the spatial connection between aggressor and victim DRAM rows, thereby offering a robust defense against complex Rowhammer access patterns, including those not mitigated by victim-focused methods like the Half-Double attack. RRS operates by periodically swapping aggressor-rows with randomly selected rows within the DRAM memory, limiting the potential damage to any single locality. While RRS can be implemented in conjunction with any tracking mechanism, its effectiveness has been demonstrated when paired with a Misra-Gries tracker, targeting a Row Hammer Threshold of 4.8K activations, akin to state-of-the-art attacks. Notably, evaluations of RRS indicate a negligible slowdown (averaging 0.4%), while providing strong security assurances against Row Hammer bit flips even under continuous attack scenarios spanning several years.
Initial beliefs held that Error Correcting Code (ECC) would serve as an effective defense against Rowhammer attacks. However, subsequent research has shown that ECC, despite its prevalence in server environments, falls short as a comprehensive solution. This inadequacy primarily arises due to ECC’s vulnerability to scenarios involving triple bit flips, a phenomenon well-documented in the literature [14]. Additionally, ECC, while standard in server-grade hardware, is typically absent in consumer-grade DRAM systems.
9.2 Adding nops To Code
A mitigation against LeapFrog attack specifically would be patching the source code or binary such that it is no longer vulnerable. Given the single bit flip requirement of Rowhammer on the PC values, adding enough nops within the LeapFrog gadget to prevent instruction skips that only require a single bit flip would potentially mitigate the attack. Adding nop instructions to source code is not trivial when the compiler optimizations are enabled since the compiler may reorder the critical parts in a different way that make the patch ineffective. A mitigation tool that adds nops to binary itself may overcome the compiler effect. Yet, adding new instructions to a binary will result in a change in the address of all the following instruction which may introduce new LeapFrog gadgets. Therefore, the patched binary needs to be re-evaluated if it still has gadgets. Although, it may potentially generate a LeapFrog proof binary, we claim it is not a sound and reliable approach.
10 Conclusion
In this work, we introduced LeapFrog, a specific type of Rowhammer exploit that directly targets the control flow of programs by manipulating the Program Counter stored in the stack. This novel approach marks a significant shift in the understanding of Rowhammer threats, moving beyond traditional data integrity attacks to those that can alter program execution. Our successful demonstration of this attack in the context of an OpenSSL TLS handshake scenario highlights its practical effectiveness and potential impact on widely used security protocols.
Furthermore, we proposed a systematic approach to identify LeapFrog gadgets in real-world software. Using our MFS analysis tool, we detected multiple points in binaries of commonly used software, such as sudo and OpenSSL’s encryption tool, which results in authentication and encryption bypass, respectively, when exploited. Even though the identification of vulnerable software is relatively straightforward thanks to our detection tool, mitigation of LeapFrog is not a trivial task since it is not transparent to the developers on a source code level. Instead, dedicated Rowhammer-resistant DRAM hardware or Rowhammer-aware compiler tools will be required to prevent LeapFrog attacks.
11 Disclaimer
Andrew Adiletta’s affiliation with The MITRE Corporation is provided for identification purposes only, and is not intended to convey or imply MITRE’s concurrence with, or support for, the positions, opinions, or viewpoints expressed by the author.
Acknowledgements
This work was supported by the National Science Foundation grant CNS-2026913 and in part by a grant from the Qatar National Research Fund.